The lengths of the sides of a triangle are consecutive even integers. Find the longest side is it is 22 units shorter than the perimeter.
Answer by ankor@dixie-net.com(22740)

The lengths of the sides of a triangle are consecutive even integers. Find the longest side if it is 22 units shorter than the perimeter.
:
Let the three sides be: x, (x+2), (x+4)
then the perimeter:
p = x + (x+2) + (x+4)
p = 3x + 6
:
"Find the longest side is it is 22 units shorter than the perimeter."
(3x+6 - (x+4) = 22
3x + 6 - x - 4 = 22
2x + 2 = 22
2x = 22 - 2
2x = 20
x = 10, the shortest side
then
10 + 4 = 14 units is the longest side
:
:
Check solution
Perimeter - longest side =
(10+12+14) - 14 = 22